# Unexpected loss of maximality: the case of Hilbert square of real surfaces

### v3

@@ 3. Cut-and-Paste construction of Hilbert squares over the reals
[…]
@@ 5.1. Reduction of the proof to computation of $\beta_{1}(X^{[2]}(\mathbb{R}))$
As before, let $\beta_{i},\,i=0,\dots,4$ denote the $\mathbb{F}_{2}$ Betti numbers of $X.$ Under assumption of Theorem 1.2, we have $\beta_{1}=\beta_{3}=0$, while $\beta_{2}$ coincides with the ordinary Betti number $b_{2}=\mbox{dim}H_{2}(X;\mathbb{Q})$.
@@ 5.2. Direct computation of $\beta_{1}(X^{[2]}(\mathbb{R}))$
For this computation, we use the cut-and-paste construction described in Section 3, and follow the notations introduced therein.
@@
$\beta_{1}(X^{[2]}_{\text{extra}}(\mathbb{R}))=$ $\sum_{i<j}(\beta_{1}(F_{i})+\beta_{1}(F_{j}))$ (5.3)
@@
and implies $\beta_{3}(F_{i}^{(2)},\Delta F_{i})=\beta_{3}(F_{i}^{(2)})+1.$ Furthermore, by applying the Macdonald formula when $F_{i}$ is oriented, and the Kallel-Salvatore formula [17] (see also [1]) in the non-orientable case, we find $\beta_{3}(F_{i}^{(2)})=\beta_{1}(F_{i}).$ Summing-up, we find $\beta_{1}(\mathbb{H}_{i})=1+\beta_{1}(F_{i})=\beta_{*}(F_{i})-1.$
@@
To finish the proof we apply the Mayer-Vietoris sequence
[…]
@@
$\operatorname{rank}\bigoplus_{i=1}^{r}\mathfrak{i}_{*}^{\mathbb{H}_{i}}=\sum^{r}_{1}\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{i}}=r,$
@@
In particular, when $X$ satisfies assumptions of Theorem 1.2, the defect of maximality of $X^{[2]}(\mathbb{R}),$ that is the difference $\beta_{*}(X^{[2]})-\beta_{*}(X^{[2]}(\mathbb{R}))$, is equal to $4(r-1).$
@@
Combining them together we get $b_{0}(X(\mathbb{R}))\geq 1+\frac{1}{2}h^{2,0}+h^{1,0}\geq 2.$
@@
[…]
@@
If $X(\mathbb{R})$ is connected (that is $r=1$), then, according to the first two items of Proposition 3.1, $\ker\mathfrak{i}_{*}^{\mathbb{H}_{0}}$ is contained in $\ker\mathfrak{i}_{*}^{\mathbb{H}_{1}}$. Therefore, $\operatorname{rank}\mu=\operatorname{rank}(\mathfrak{i}_{*}^{\mathbb{H}_{0}}\oplus\mathfrak{i}_{*}^{\mathbb{H}_{1}})=\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}},$ while by (6.3) and Proposition 3.1.3) we have $\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=\beta_{1}(\mathbb{H}_{0})=1+\beta_{1}$.
Now, assume that $r>1+\beta_{1}$ and show, first, that under this assumption the homomorphism $\mathfrak{i}_{*}^{\mathbb{H}_{0}}:H_{1}(\bigsqcup_{i=1}^{r}\mathbb{P}_{\mathbb{R}}T^{*}F_{i})\to H_{1}(\mathbb{H}_{0})$ vanishes on at least one nontrivial linear combination of the fiber-classes $w_{i},i=1,\dots,r$.
Pick a point $p_{i}$ on each $F_{i}\subset X$ with $i=1,\dots,r$. For each $i=2\dots,r$, join $p_{1}$ with $p_{i}$ by a generic path $\gamma_{i}$ in $X$ and consider the circles $C_{i}$ formed by $\gamma_{i}\cup\operatorname{c}(\gamma_{i})$. Since their number is strictly bigger than $\beta_{1}(X)$, there exists a 2-dimensional submanifold $\Upsilon\subset X$ bounding some nontrivial combination of circles $C_{i}$. The projection $\operatorname{pr}(\Upsilon)$ of $\Upsilon$ into $X/\operatorname{c}$ gives a 2-cycle which has a nontrivial $\mathbb{F}_{2}$-intersection with $F_{i}\subset X/\operatorname{c}$ for at least one of $i=2\dots,r$. This implies that at least one of $F_{i}$ represents a nonzero element of $H_{2}(X/\operatorname{c})$.
@@
it is sufficient now to pick an $F_{k}$ representing a nonzero element in $H_{2}(X/\operatorname{c})$ and to take a generic smooth $2$-dimensional submanifold $\Sigma\subset X/\operatorname{c}$ which has nontrivial $\mathbb{F}_{2}$-intersection with this $F_{k}.$ Then, the linear combination of the fiber-classes $w_{i}$ for which $F_{i}$ has a non-zero $\mathbb{F}_{2}$-intersection with $\Sigma$ vanishes.
According to the first two items of Proposition 3.1, any nontrivial combination of fiber-classes is not mapped to zero by $\bigoplus_{i=1}^{r}\mathfrak{i}_{*}^{\mathbb{H}_{i}}.$ Therefore, we conclude that $\operatorname{rank}\mu=\operatorname{rank}(\mathfrak{i}_{*}^{\mathbb{H}_{0}}\bigoplus_{i=1}^{r}\mathfrak{i}_{*}^{\mathbb{H}_{i}})>\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}$. Finally, there remains to notice that by (6.3) and the last item of Proposition 3.1 we have $\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=1+\beta_{1}.$
@@ 7.1.3. Abelian surfaces
No real abelian surfaces has a maximal Hilbert square.
@@
This shows that the sum of the 4 fiber-classes $w_{k}\in H_{1}(P_{\mathbb{R}}T^{*}F_{k})$ belongs to $\ker\mathfrak{i}_{*}^{\mathbb{H}_{0}}$ and that $\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=1+\beta_{1}$. On the other hand, by Proposition 3.1, neither of nontrivial combinations of $w_{k}$ is mapped to 00 by $\bigoplus_{i=1}^{4}\mathfrak{i}_{*}^{\mathbb{H}_{i}}$. Therefore, $\operatorname{rank}\mu=\operatorname{rank}(\mathfrak{i}_{*}^{\mathbb{H}_{0}}\bigoplus_{i=1}^{4}\mathfrak{i}_{*}^{\mathbb{H}_{i}})>\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=1+\beta_{1}.$
For comparison with the case of ruled surfaces, let us note that if $X$ is a maximal abelian surface, we have $41<\beta_{0}(X(\mathbb{R}))=\beta_{1}(X)=$ 4 .
@@ 7.1.4. Rational surfaces
The Hilbert square $X^{[2}]$ is not maximal if $X$ is a real del Pezzo surface with $K_{X}^{2}=1$ and $X(\mathbb{R})$ homeomorphic to $\mathbb{P}^{2}(\mathbb{R})\sqcup 4S^{2}$, or its blowup at some number of real points. For all other maximal real rational surfaces $X$, their Hilbert square is maximal.
@@ 7.1.5. Product of curves
According to Theorem 1.5, if $X$ is a maximal real surface such that $b_{0}(X(\mathbb{R}))>1+\beta_{1},$ its Hilbert square is not maximal. Examples of such surfaces with $\beta_{1}>0$ exist in abundance. The simplest ones are given by products $C_{1}\times C_{2},$ where $C_{1}$ and $C_{2}$ are maximal curves of genus $g_{1}$ and $g_{2},$ respectively, where $g_{1},g_{2}\geq 2$ and $g_{1}+g_{2}>4.$
@@ 7.1.6. Enriques surfaces
For surfaces $X$ with $\operatorname{Tors}_{2}H^{2}(X;\mathbb{Z})\neq 0,$ one can still compute the $\mathbb{F}_{2}$-Betti numbers of the Hilbert square $X^{[2]}$ for a given example, albeit a general formula is lacking [28, Example 2.5]. We address here the maximality of the Hilbert square for real Enriques surfaces, and we find:
@@
No real Enriques surfaces has a maximal Hilbert square.
@@
As mentioned in Introduction, the real locus of a maximal real Enriques surface is always disconnected. This can be proved directly, i.e., without appealing to a full classification [7], in the following manner. Namely, let us assume that the real locus $X(\mathbb{R})$ of some maximal real Enriques surface $X$ is connected. Then, $X(\mathbb{R})$ is homeomorphic to a real projective plane blown-up at $16-3=13$ real points. Its universal covering is a double covering $Y\to X,$ where $Y$ is a K3 surface. $Y$ can be equipped with two possible lifts of the real structure. We choose the one for which $Y(\mathbb{R})\neq\emptyset.$ Then, $Y(\mathbb{R})$ is the orientation double cover of $X(\mathbb{R}),$ and so $\chi(Y(\mathbb{R}))=2\chi(X(\mathbb{R}))=-24.$ Since $Y(\mathbb{R})$ is orientable, it follows that $Y(\mathbb{R})$ is homeomorphic to a sphere with 13 handles, which contradicts the Smith inequality $\beta_{*}(Y(\mathbb{R}))\leq\beta_{*}(Y)=24$.
@@ 7.2. Application to cubic 4-folds
The deformation classification of real nonsingular cubic 4-folds, established by S. Finashin and V. Kharlamov [9], associates with all but one class (called by them irregular) a deformation class of a real nonsingular K3 surface. On the other hand, due to V. Krasnov [22], for each regular class of cubic 4-folds (in Finashin-Kharlamov’s sense) except the one which corresponds to a K3 with 10 spheres as the real locus, the Fano variety is equivariantly diffeomorphic to the Hilbert square of the corresponding K3. In a separate paper [23], Krasnov proved that for the cubics in the irregular class, the real locus of the Fano variety is a disjoint union of six disjoint copies of $S^{2}\times S^{2}$ and one more component which is homeomorphic to $X^{[2]}_{\text{main}}(\mathbb{R}),$ where $X$ is a K3-surface with $X(\mathbb{R})=3S^{2}$. Therefore, combining these results with Corollary 1.3 and the computations from the proof of Theorem 1.2, we get the following statement.

### v4

@@ 3. Cut-and-Paste construction of Hilbert squares over the reals
${H_{G}^{*}(X^{[2]})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R^{[2]}}$${H_{G}^{*}(\operatorname{Bl}_{p}(X))\,\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R}$${H^{*}(X^{[2]})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}}$${H^{*}(\operatorname{Bl}_{p}(X)).}$
@@ 5.1. Reduction of the proof to computation of $\beta_{1}(X^{[2]}(\mathbb{R}))$
As before, let $\beta_{i},\,i=0,\dots,4$ denote the $\mathbb{F}_{2}$ - Betti numbers of $X.$ Under assumption of Theorem 1.2, we have $\beta_{1}=\beta_{3}=0$, while $\beta_{2}$ coincides with the ordinary Betti number $b_{2}=\mbox{dim}H_{2}(X;\mathbb{Q})$.
@@ 5.2. Direct computation of $\beta_{1}(X^{[2]}(\mathbb{R}))$
The cut-and-paste construction described in Section 3 is used next to compute $\beta_{1}(X^{[2]}(\mathbb{R})).$ We follow the notations introduced therein.
@@
[…]
@@
and implies $\beta_{3}(F_{i}^{(2)},\Delta F_{i})=\beta_{3}(F_{i}^{(2)})+1.$ Furthermore, by applying the Macdonald formula [28, Proposition 4.2] expressing the Betti numbers of symmetric products when $F_{i}$ is oriented, and the Kallel-Salvatore formula [18, Proposition 20] (see [1, Theorem 3.3] for a more precise formulation) in the non-orientable case, we find $\beta_{3}(F_{i}^{(2)})=\beta_{1}(F_{i}).$ Summing-up, we find $\beta_{1}(\mathbb{H}_{i})=1+\beta_{1}(F_{i})=\beta_{*}(F_{i})-1.$
@@
To finish the proof we apply the Mayer-Vietoris sequence to the decomposition $X^{[2]}=A\cup B$ with $A=\mathbb{H}_{0}$ and $B=\sqcup_{i=1}^{r}\mathbb{H}_{i}:$
$\bigoplus_{i=1}^{r}H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{i})\xrightarrow{\mu}\bigoplus_{i=0}^{r}H_{1}(\mathbb{H}_{i})\rightarrow H_{1}(X^{[2]}_{\text{main}}(\mathbb{R}))\rightarrow 0$
@@
$\operatorname{rank}\bigoplus_{i=1}^{r}\mathfrak{i}_{*}^{\mathbb{H}_{i}}=\sum^{r}_{i=1}\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{i}}=r,$
@@
In particular, when $X$ satisfies assumptions of Theorem 1.2, the defect of maximality of $X^{[2]}(\mathbb{R}),$ that is the difference $\beta_{*}(X^{[2]})-\beta_{*}(X^{[2]}(\mathbb{R}))$, is equal to 4(r1).4(r-1).
@@
Combining them together we get $b_{0}(X(\mathbb{R}))\geq 1+\frac{1}{2}h^{2,0}+h^{1,0},$ which implies $b_{0}(X(\mathbb{R}))\geq 2.$
@@
$\bigoplus_{i=1}^{r}H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{i})\xrightarrow{\mu}\bigoplus_{i=0}^{r}H_{1}(\mathbb{H}_{i})\rightarrow H_{1}(X^{[2]}_{\text{main}}(\mathbb{R}))\rightarrow 0.$ (6.1)
@@
If $X(\mathbb{R})$ is connected (that is $r=1$), then, according to Corollary 3.2, $\ker\mathfrak{i}_{*}^{\mathbb{H}_{0}}$ is contained in $\ker\mathfrak{i}_{*}^{\mathbb{H}_{1}}$. Therefore, $\operatorname{rank}\mu=\operatorname{rank}(\mathfrak{i}_{*}^{\mathbb{H}_{0}}\oplus\mathfrak{i}_{*}^{\mathbb{H}_{1}})=\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}},$ while by (6.3) and Proposition 3.1.(3) we have $\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=\beta_{1}(\mathbb{H}_{0})=1+\beta_{1}$.
Assume now that $r>1+\beta_{1}.$ Under this assumption, the homomorphism, we show first that $\mathfrak{i}_{*}^{\mathbb{H}_{0}}:H_{1}(\bigsqcup_{i=1}^{r}\mathbb{P}_{\mathbb{R}}T^{*}F_{i})\to H_{1}(\mathbb{H}_{0})$ vanishes on at least one non-trivial linear combination of the fiber-classes $w_{i},\,i=1,\dots,r.$
Pick a point $p_{i}$ on each $F_{i}\subset X$ with $i=1,\dots,r$. For each $i=2\dots,r,$ join $p_{1}$ with $p_{i}$ by a generic path $\gamma_{i}$ in $X$ and consider the circles $C_{i}$ formed by $\gamma_{i}\cup\operatorname{c}(\gamma_{i}).$ Since their number, r1,r-1, is strictly larger than $\beta_{1}(X),$ there exists a 2-dimensional submanifold $\Upsilon\subset X$ bounding some nontrivial combination of circles $C_{i}.$ For convenience, we choose $\Upsilon$ such that it is transversal to $\sqcup_{i=1}^{r}F_{i}.$ Its image $\operatorname{pr}\Upsilon$ under the projection map $\operatorname{pr}:X\rightarrow X/\operatorname{c}$ is a $2$-cycle, giving rise to an element $\operatorname{pr}\Upsilon]\in H_{2}(X/\operatorname{c}).$ For each $i$ such that $C_{i}$ is a part of $\partial\Upsilon,$ let $q\in F_{i}\cap\Upsilon$ and compute the local intersection number of $\operatorname{pr}\Upsilon$ and $F_{i}$ at $\operatorname{pr}(q)$ in $X/\operatorname{c}.$ By the projection formula, we have
$\left([\operatorname{pr}\Upsilon]\circ F_{i}\right)_{\operatorname{pr}(q)}=$ $\frac{1}{2}\left(\operatorname{pr}^{*}[\operatorname{pr}\Upsilon]\circ\operatorname{pr}^{*}F_{i}\right)_{q}$
$=$ $\left(\operatorname{pr}^{*}[\operatorname{pr}\Upsilon]\circ F_{i}\right)_{q}$
$=$ $\left([\Upsilon\cup\operatorname{c}(\Upsilon)]\circ F_{i}\right)_{q}.$ (6.4)
If $q=p_{i},$ the cycle $\Upsilon\cup\operatorname{c}(\Upsilon)$ intersects $F_{i}$ transversally at the point $p_{i}.$ By (6), we find that the local intersection number $\left([\operatorname{pr}\Upsilon]\circ F_{i}\right)_{\operatorname{pr}(q)}$ equals to $1.$ If $q\neq p_{i},$ then each $\Upsilon$ and $\operatorname{c}(\Upsilon)$ intersects $F_{i}$ transversally at $q$, and so $([\Upsilon\cup\operatorname{c}(\Upsilon)]\circ F_{i})_{q}=2$. We conclude that the cycle $\operatorname{pr}\Upsilon$ has a non-trivial $\mathbb{F}_{2}$-intersection with $F_{i}\subset X/\operatorname{c}$ for some $i\in\{2\dots,r\}.$ This implies that there exists at least one $i\in\{2,\dots,r\}$ such that $F_{i}$ represents a non-zero element of $H_{2}(X/\operatorname{c}).$
@@
we identify $\bigsqcup_{i=1}^{r}\mathbb{P}_{\mathbb{R}}T^{*}F_{i}$ with the boundary of a tubular neighborhood $U$ of $\cup_{i=1}^{r}F_{i}$ in $X/\operatorname{c}$, and $\mathbb{H}_{0}$ with $X/\operatorname{c}\setminus\operatorname{Int}U$. Next, we pick $F_{k}$ representing a non-zero element in $H_{2}(X/\operatorname{c})$ and take a generic smooth $2$-dimensional submanifold $\Sigma\subset X/\operatorname{c}$ which has non-trivial $\mathbb{F}_{2}$-intersection with this $F_{k}.$ Then, the linear combination of the fiber classes given by $\Sigma\cap\partial U\subset\partial U=\bigsqcup_{i=1}^{r}\mathbb{P}_{\mathbb{R}}T^{*}F_{i}$ is bounded by $\Sigma\setminus\operatorname{Int}U$ and, thus, vanishes in $H_{1}(\mathbb{H}_{0})$.
According to the first item of Proposition 3.1, any non-trivial combination of fiber-classes is not mapped to zero by $\bigoplus_{i=1}^{r}\mathfrak{i}_{*}^{\mathbb{H}_{i}}.$ Therefore, we conclude that $\operatorname{rank}\mu=\operatorname{rank}(\mathfrak{i}_{*}^{\mathbb{H}_{0}}\bigoplus_{i=1}^{r}\mathfrak{i}_{*}^{\mathbb{H}_{i}})>\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}.$ Finally, there remains to notice that, by (6.3) and the last item of Proposition 3.1, we have $\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=1+\beta_{1}.$
@@ 7.1.3. Abelian surfaces
No real abelian surfaces has maximal Hilbert square.
@@
This shows that the sum of the 4 fiber-classes $w_{k}\in H_{1}(P_{\mathbb{R}}T^{*}F_{k})$ belongs to $\ker\mathfrak{i}_{*}^{\mathbb{H}_{0}}$ and that $\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=1+\beta_{1}$. On the other hand, by Proposition 3.1, neither of nontrivial combinations of $w_{k}$ is mapped to 0 by $\bigoplus_{i=1}^{4}\mathfrak{i}_{*}^{\mathbb{H}_{i}}$. Therefore, $\operatorname{rank}\mu=\operatorname{rank}(\mathfrak{i}_{*}^{\mathbb{H}_{0}}\bigoplus_{i=1}^{4}\mathfrak{i}_{*}^{\mathbb{H}_{i}})>\operatorname{rank}\mathfrak{i}_{*}^{\mathbb{H}_{0}}=1+\beta_{1}.$
For comparison with the case of ruled surfaces, let us note that if $X$ is a maximal abelian surface, we have $1<\beta_{0}(X(\mathbb{R}))=\beta_{1}(X)=4.$
@@ 7.1.4. Rational surfaces
The Hilbert square $X^{[2]}$ is not maximal if $X$ is a real del Pezzo surface with $K_{X}^{2}=1$ and $X(\mathbb{R})$ homeomorphic to $\mathbb{P}^{2}(\mathbb{R})\sqcup 4S^{2}$, or its blowup at some number of real points. For all other maximal real rational surfaces $X$, their Hilbert square is maximal.
@@ 7.1.5. Product of curves
According to Theorem 1.5, if $X$ is a maximal real surface such that $b_{0}(X(\mathbb{R}))>1+\beta_{1},$ its Hilbert square is not maximal. Examples of such surfaces with $\beta_{1}>0$ exist in abundance. The simplest ones are the products $C_{1}\times C_{2},$ where $C_{1}$ and $C_{2}$ are maximal curves of genus $g_{1}$ and $g_{2},$ respectively, where $g_{1},g_{2}\geq 2$ and $g_{1}+g_{2}>4.$
@@ 7.1.6. Enriques surfaces
For surfaces $X$ with $\operatorname{Tors}_{2}H^{2}(X;\mathbb{Z})\neq 0,$ one can still compute the $\mathbb{F}_{2}$-Betti numbers of the Hilbert square $X^{[2]}$ for a given example, albeit a general formula is lacking [31, Example 2.5]. We address here the maximality of the Hilbert square for real Enriques surfaces, and we find:
@@
No real Enriques surfaces has maximal Hilbert square.
@@
As mentioned in Introduction, the real locus of a maximal real Enriques surface is always disconnected. This can be proved directly, i.e., without appealing to a full classification [7], in the following manner. Namely, let us assume that the real locus $X(\mathbb{R})$ of some maximal real Enriques surface $X$ is connected. Then, $X(\mathbb{R})$ is homeomorphic to a real projective plane blown-up at 163=1316-3=13 real points. Its universal covering is a double covering $Y\to X,$ where $Y$ is a K3 surface. $Y$ can be equipped with two possible lifts of the real structure. We choose the one for which $Y(\mathbb{R})\neq\emptyset.$ Then, $Y(\mathbb{R})$ is the orientation double cover of $X(\mathbb{R}),$ and so $\chi(Y(\mathbb{R}))=2\chi(X(\mathbb{R}))=-24.$ Since $Y(\mathbb{R})$ is orientable, it follows that $Y(\mathbb{R})$ is homeomorphic to a sphere with 13 handles, which contradicts the Smith inequality $\beta_{*}(Y(\mathbb{R}))\leq\beta_{*}(Y)=24$.
@@ 7.2. Application to cubic 4-folds
[…]